OpenStudy (anonymous):

Solve –2(x – 6) = –4 Solve x – 2y = 4 for y.

OpenStudy (anonymous):

Are they wanting to use x from the previous problem?

OpenStudy (anonymous):

no i ment to put that they are two difffrent Q

OpenStudy (anonymous):

if so that would be 8 - 2y = 4 2y - 8 = 4 -8 -8 2y = -4 y = -2 if not it would be 1 - 2y = 4 2y - 1 = 4 2y = 3 y = .6 with a repeating bar above the 6

OpenStudy (anonymous):

x automatically is substituted for 1 because it's simply x, so that's how I got the 1, if you were wondering.